OpenStudy (anonymous):

How would you simplify the following expression: (xy^6/4x^3y)^-2

OpenStudy (anonymous):

flip it upside down and then square everything. remember only to add powers and square coefecients.

OpenStudy (anonymous):

\[(xy^6/4x^3y)^-2\] \[(4x^33y/xy^6)^2\] \[144x^6y^2/x^2y^(12)=144x^4/y^(10)\]

OpenStudy (anonymous):

a^(-2) = 1/ a^2

OpenStudy (anonymous):

how did you come up with 144?

OpenStudy (anonymous):

read the power as a coefficent it should be 16 not 144 im sorry

OpenStudy (anonymous):

so the second eq. should not have a 3y

OpenStudy (anonymous):

Still not understanding, but ty for trying

OpenStudy (anonymous):

ok so take the original eq and flip it to get rid of the ^-2 to a ^2

OpenStudy (anonymous):

then it looks like this \[(xy^6/4x^3y)^-2=(4x^3y)^2/(xy^6)^2\]

OpenStudy (anonymous):

then square the top and the bottom 16x^6y^2/x^2y^12

OpenStudy (anonymous):

\[16x^6y^2/x^2y^12=16(x^4/y^10)\]

OpenStudy (anonymous):

wouldn't the y^6 be y^36?

OpenStudy (anonymous):

because you don't multiply it by 2 rather 6 x 6?

OpenStudy (anonymous):

no remeber (x^2)^4does not equalx^16 x^2x^2x^2x^2=(x^2)^4 so what you do is add the powers

OpenStudy (anonymous):

think of(x^2)^4 as 2*4=8 you simply multiply a power to a power

OpenStudy (anonymous):

ok, can you explain how you squared the first equation then, because the only part I'm getting is the 16 and the y^2

OpenStudy (anonymous):

but you understand why you flip it right?

OpenStudy (anonymous):

\[(4x^3y)^2=(4^2(x^3)^2y^2)\]

OpenStudy (anonymous):

(4x^3y)^2, I get how you got the 16, 4^2 = 16, but would the (x^3)^2 = x^6 or x^5?

OpenStudy (anonymous):

\[(4x^3y)^2=((4x^3y))((4x^3y))\]

OpenStudy (anonymous):

and then use FOIL?

OpenStudy (anonymous):

you either think of (x^3)^2 as x^3x^3=x^6 remeber you can add the powers or if its a power to a power you multiply 3*2=6

OpenStudy (anonymous):

^^ ryan its y^10 not y^8

OpenStudy (anonymous):

im sorry y^12

OpenStudy (anonymous):

thank you, I'm sorry I'm such an idiot about this stuff, I don't get how y'all can figure this stuff out. I have 20 math problems to get done by midnight and I'm only on #5.................... I don't think I'll get this done....lol